# Moving toward best practice when using propensity score weighting in survey observational studies

### v1

@@ 6 Discussion
Our paper introduces a unified framework for incorporating survey weights into the balancing weights paradigm, enhancing their applicability to survey observational data and providing robust tools for population-level causal inference. We establish the asymptotic normality of survey-weighted estimators for retrospective designs, with extensions to prospective cases detailed in the appendix. A robust sandwich variance estimator for weighting methods is developed via M-estimation theory, which extends to a unified variance framework for all three augmented estimators— one-step moment estimator (MOM), clever covariate regression estimator (CVR), and weighted regression estimator (WET)—derived from the shared augmented weighting estimating equations for survey observational data. Simulation studies evaluate these methods under varying overlap and model misspecification conditions. Results demonstrate WET’s robustness, MOM’s consistency, and CVR’s efficiency. Practical applicability is illustrated through two data examples: estimating the effect of special education services on fifth-grade math achievement using the Early Childhood Longitudinal Study Kindergarten Class (ECLS-K) dataset under a prospective multi-stage sampling survey design, and analyzing racial disparities in healthcare expenditures using the Medical Expenditure Panel Survey (MEPS) under a retrospective descriptive comparison framework. These case studies highlight the strengths of survey-weighted propensity score models (W.PS) in achieving population-level covariate balance, WET in providing robust estimation, and the PATO estimand in enabling reliable inference within comparable subpopulations. All proposed methods are implemented in the publicly available PSweight R package.
First, incorporating survey weights throughout all phases of the propensity score weighting estimator is essential for unbiased treatment effects estimation in survey observational studies. Using a survey-weighted regression model (W.PS) for propensity score estimation aligns sampled data covariate distributions with the target population, enhancing efficiency, coverage, and robustness across varying overlap scenarios. Consistent with Ridgeway et al.,[4] our results suggest that omitting survey weights restricts balance to within the sample,[2, 46] potentially introducing bias when generalizing to target populations. Consequently, W.PS is integral to valid population-level causal inferences.
@@
Second, augmented estimators are instrumental in reducing bias, enhancing efficiency, and improving coverage in survey observational data, particularly when both the propensity score and outcome models are correctly specified. Simulation results hightlight the strengths of MOM, CVR and WET in different scenarios. WET demonstrates superior robustness under model misspecification and consistently performs well across estimands and overlap conditions. CVR shows high efficiency under good overlap and correct models but exhibits sensitivity to outcome model misspecification and diminished reliability in poor overlap scenarios, limiting its utility in real-world applications. These findings align with prior literature [30, 26] that highlights the instability of inverse propensity-based covariate approaches, particularly under extreme values. Extrapolation issues are exacerbated when the clever covariate incorporates large survey weights, amplifying variability and undermining model performance. Reformulating the clever covariate to preserve boundedness properties[26] could enhance its robustness and practical utility, particularly in scenarios with extreme survey weights or limited overlap.
@@
Third, WET consistently emerges as the most robust and efficient estimator across diverse overlap conditions and model misspecifications. Compared to CVR and MOM, WET demonstrates superior stability and precision, particularly in poor overlap or when models are misspecified. This aligns with Kang and Schafer[30] and Robins et al.[26], who suggest that clever covariates can supplement outcome models but fail to address misspecifications under poor overlap, leading to coverage declines as low as 30%, as shown in Table 6. Moreover, as noted by Robins et al.,[26] when the CVR is structured with an identity link, it aligns with Targeted Maximum Likelihood Estimation (TMLE) for marginal means but can perform worse than the weighted regression estimator (WET) when both models are misspecified. Simulation results and case studies further confirm WET’s robustness, showing lower bias and more reliable coverage in both prospective and retrospective survey designs. Case study analyses demonstrate WET’s capacity to mitigate the influence of extreme survey weights, particularly for PATO, deliving stable and interpretable results. Overall, WET under W.PS is recommended for complex survey settings, combining efficiency, robustness, and practical reliability.
@@
The validity of our study’s findings hinges on critical assumptions, including the stable unit treatment value assumption (SUTVA) and treatment assignment ignorability. [47, 48] These assumptions are crucial for unbiased causal effect estimation under observational designs but may be violated in real-world settings due to issues such as latent confounders or non-overlapping covariate distributions.[49, 50] Hence, sensitivity analyses could provide critical insights into the potential impact of unmeasured confounding for weighting and augmented weighting estimator, and need to be generalized to complex survey data.[51] Second, our analysis assumes that survey weights are known and derived from well-designed surveys. If this assumption does not hold—for instance, in cases where the survey design introduces biases or weights are not appropriately generated—causal inferences may be distorted. Poor survey design, including unaddressed non-response biases or inadequate post-stratification, can result in flawed survey weights that exacerbate selection bias or produce extreme values, destabilizing estimators and reducing efficiency.[50, 19] Future research could extend sensitivity analyses to explore the robustness of survey weights under alternative assumptions about the sampling mechanism.[49, 52] Third, while our findings focus on binary treatments, they are also readily extensible to multiple treatments. With multiple nominal or ordinal treatments, the generalized propensity scores (GPS) provide a key device for simultaneously balancing covariates across multiple treatment arms, [53, 45] and are key ingredients in developing weighting and augmented weighting estimators. Future work is needed to extend the current methods to multiple treatments in the context of survey data. Finally, our analysis used parametric models for propensity score and outcome estimation, as these are the most commonly used methods in real practice. However, these parametric models may fail to capture complex, high-dimensional, or nonlinear structures in real-world data [54]. More flexible approaches, such as Bayesian Additive Regression Trees (BART) [55, 56] and Causal Forests [57] can effectively capture non-linear relationships and heterogeneous treatment effects. These flexible modeling frameworks are particularly advantageous in high-dimensional settings or when traditional parametric models prove inadequate.[58] Advanced estimators like the Targeted Maximum Likelihood Estimation (TMLE) can achieve semiparametric efficiency by incorporating machine learning methods to optimize both outcome and propensity score estimates.[59] Extensions of AIPW methods to double machine learning (DML) frameworks can also further enhance robustness by addressing model misspecification and overfitting challenges.[54] These data-adaptive extensions of the current estimators in the context of survey observational data represent fruitful directions for additional research.

### v2

@@ 6 Discussion
We introduce a unified framework for incorporating survey weights into the balancing weights estimators for population-level causal inference. We establish the asymptotic normality of survey-weighted estimators for retrospective designs (with prospective designs detailed in the Web Appendix), and provide a robust sandwich variance estimator for weighting and augmented estimators. Simulation studies evaluate these methods under varying overlap and model misspecification conditions, and highlight weighted regression estimator as a numerically stable approach even in challenging poor overlap settings. These case studies highlight the strengths of survey-weighted propensity score models (W.PS) in achieving population-level covariate balance, WET in providing robust estimation, and the PATO estimand in enabling reliable inference within comparable subpopulations. All proposed methods are implemented in the publicly available PSweight R package.
Several lessons have emerged from our exploration. First, incorporating survey weights throughout all phases of the propensity score weighting estimator is essential for unbiased treatment effects estimation in survey observational studies. Using a survey-weighted regression model (W.PS) for propensity score estimation aligns sampled data covariate distributions with the target population, enhancing efficiency, coverage, and robustness across varying overlap scenarios. Consistent with Ridgeway et al.,[4] our results suggest that omitting survey weights restricts balance to within the sample,[2, 41] potentially introducing bias when generalizing to target populations. Consequently, W.PS is integral to valid population-level causal inferences. Our investigation clarifies when survey-weighted propensity scores are necessary versus beneficial. Under retrospective (informative) sampling where $e_{\text{fp}}(\mathbf{x})\neq e_{\text{sp}}(\mathbf{x})$, Equation (1) shows that survey-weighted estimation is essential for recovering the population propensity score required by the balancing weights. Under prospective (non-informative) sampling where these quantities coincide, either approach is theoretically valid. However, our simulations (Tables 5-6) demonstrate consistent practical advantages of survey-weighted estimation (W.PS) over alternatives across all scenarios. We therefore recommend using survey-weighted propensity score models as the default when survey weights are available, particularly for weighting methods targeting population effects. An important practical consideration concerns the specification of $e_{\text{fp}}(\mathbf{x})$ under retrospective sampling. While Equation (1) reveals a complex nonlinear form, our sensitivity analyses (Web Appendix B.3) demonstrate that under rare sampling conditions (survey weights >10>10, corresponding to sampling rates $10\%$), a standard unweighted logistic regression provides a robust and computationally feasible approximation for $e_{\text{fp}}(\mathbf{x})$ when combined with survey-weighted estimation of $e_{\text{sp}}(\mathbf{x})$ in augmented estimators. For studies with higher sampling rates (survey weights <10<10), researchers should consider explicitly modeling the nonlinear form through maximum likelihood estimation that incorporates knowledge of the sampling mechanism, though this requires stronger assumptions about the sampling design and increased computational complexity.
@@
Second, augmented estimators are instrumental in reducing bias, enhancing efficiency, and improving coverage in survey observational data, particularly when both the propensity score and outcome models are correctly specified. Our extensive simulation results highlight the strengths of MOM, CVR and WET in different scenarios. WET demonstrates superior robustness under model misspecification and consistently performs well across estimands and overlap conditions. CVR shows high efficiency under good overlap and correct models but exhibits sensitivity to outcome model misspecification and diminished reliability in poor overlap scenarios. These findings align with prior literature [26, 27] that highlights the instability of inverse propensity-based covariate approaches (without survey sampling), particularly under extreme values. Extrapolation issues are exacerbated when the clever covariate incorporates large survey weights, amplifying variability and undermining model performance. Reformulating the clever covariate to preserve boundedness properties[27] could enhance its robustness and practical utility, particularly in scenarios with extreme survey weights or limited overlap, and is an area for future research.
@@
Third, WET consistently emerges as the most robust and efficient estimator across diverse overlap conditions and model misspecifications. Compared to CVR and MOM, WET demonstrates superior stability and precision, particularly in poor overlap or when models are misspecified. This aligns with Kang and Schafer[26] and Robins et al.[27], who suggest that clever covariates can supplement outcome models but fail to address misspecifications under poor overlap, leading to coverage declines as low as 30%, as shown in Table 6. Moreover, as noted by Robins et al.,[27] when the CVR is structured with an identity link, it aligns with Targeted Maximum Likelihood Estimation (TMLE) for marginal means but can perform worse than the weighted regression estimator (WET) when both models are misspecified. Simulation results and case studies further confirm the robustness of WET, showing lower bias and more reliable coverage in both prospective and retrospective survey designs. Our case study analyses also demonstrate the capacity of WET to mitigate the influence of extreme survey weights, particularly for PATO. Overall, WET under W.PS can be considered as a recommended approach for causal inference with complex surveys.
@@
The validity of our study’s findings hinges on critical assumptions, including the stable unit treatment value assumption (SUTVA) and treatment assignment ignorability. [42] These assumptions are crucial for unbiased causal effect estimation under observational designs but may be violated. Hence, sensitivity analyses [43] could provide critical insights into the potential impact of unmeasured confounding for weighting and augmented weighting estimator, and need to be generalized to complex survey data. Second, our analysis assumes that survey weights are known and derived from well-designed surveys. If this assumption fails to hold, causal inferences may be compromised. Insufficient survey designs, including unaddressed non-response biases or inadequate post-stratification, can result in flawed survey weights that exacerbate selection bias or produce extreme values, destabilizing estimators and reducing efficiency. Future research could extend sensitivity analyses to explore the robustness of survey weights under alternative assumptions about the sampling mechanism. When design weights are unavailable entirely, as in voluntary or nonprobability samples, alternative approaches such as generalized entropy calibration[44] offer complementary frameworks for selection bias adjustment. Third, while our findings focus on binary treatments, they are also readily extensible to multiple treatments. With multiple nominal or ordinal treatments, the generalized propensity scores (GPS) provide a key device for simultaneously balancing covariates across multiple treatment arms, [45] and are key ingredients in developing weighting and augmented weighting estimators. Future work is needed to extend the current methods to multiple treatments in the context of survey data. Finally, our analysis relied on parametric models for propensity score and outcome estimation, reflecting their widespread use in applied practice. Nevertheless, data-adaptive machine learning approaches offer greater flexibility in capturing nonlinear relationships and complex interactions, and may be particularly advantageous in high-dimensional settings or when parametric assumptions are violated. Extending the proposed AIPW framework to incorporate double machine learning estimators based on data-adaptive working models [46] for survey-weighted observational data represents an important and promising direction for future research.